How quickly can you do this? Fill appropriate sign. ( ‘’) 
2 / 4 ☐ 3 / 6

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

2 / 4 = 1 / 2 and 3 / 6 = 1 / 2
∴ 2 / 4 = 3 / 6

Thanks